OpenStudy (anonymous):

What is the 8th term of the geometric sequence 1, –4, 16, …? Answer 4,096 16,384 –4,096 –16,384

OpenStudy (ash2326):

a=1 r=-4 an=a*r^(n-1) a8=1*-4^(7) a8=-16384

OpenStudy (anonymous):

Thank you :D

OpenStudy (ash2326):

you're welcome, don't have to say thanks